OpenStudy (liv1234):

In one generation, the frequency of the A allele is 70 percent and the frequency of the a allele is 30 percent. What are the chances of an individual in the next generation having the genotype aa if the population is in genetic equilibrium? A. 49% B. 42% C. 30% D. 9% I will give medal for best answer thank you!(:

OpenStudy (liv1234):

@wolfe8 @TheSmartOne can you guys help me out please?

OpenStudy (liv1234):

@sleepyjess @EclipsedStar

OpenStudy (aaronq):

f(a)= 0.3 f(aa)=\(a^2=(0.3)^2\) recall: \(a^2+2ab+b^2=1\)

OpenStudy (liv1234):

That doesn't make sense for this question

OpenStudy (liv1234):

@aaronq

OpenStudy (aaronq):

\(\sf f(aa)=a^2=(0.3)^2=0.09=9\%\) have you been paying attention in class?

OpenStudy (liv1234):

Yes, I have been. I just needed an explanation thank you!

OpenStudy (aaronq):

np!